# Narrow line AGN selection in CEERS: spectroscopic selection, physical properties, X-ray and radio analysis

### v1

@@
Narrow line AGN selection in CEERS: spectroscopic selection, physical properties, X-ray and radio analysis
The transformative era opened by JWST on the high- Universe allows us to investigate the early stages of the Supermassive Black Holes (SMBH) evolution, with the first results showing a greater than expected number of active galactic nuclei (AGN) at very early times. In this work, we spectroscopically select narrow-line AGN (NLAGN) among the publicly available medium-resolution spectra of the Cosmic Evolution Early Release Science Survey (CEERS). Using both traditional and newly identified emission line NLAGN diagnostics diagrams, we identified 52 NLAGN at on which we performed a detailed multiwavelength analysis. We also identified 4 new broad-line AGN (BLAGN), in addition to the 8 previously reported high-redshift BLAGN. We found that the traditional BPT diagnostic diagrams are not suited to identify high- AGN, while most of the high- NLAGN are selected using the recently proposed AGN diagnostic diagrams based on the [O iii] 4363 auroral line or high-ionization emission lines.We compared the emission line velocity dispersion and the obscuration levels of the sample of NLAGN with those of the parent sample without finding significant differences between the two distributions, suggesting a population of AGN heavily buried and not significantly impacting the host galaxies’ physical properties, as further confirmed by SED-fitting. The bolometric luminosities of the high- NLAGNs selected in this work are well below those sampled by surveys before JWST, potentially explaining the weak impact of these AGN. Finally, we investigate the X-ray properties of the selected NLAGN and of the sample of high- BLAGN. We find that all but 4 NLAGN are undetected in the deep X-ray image of the field, as well as all the high- BLAGN. We do not obtain a detection even by stacking the undetected sources, resulting in an X-ray weakness of dex from what is expected based on their bolometric luminosities. To discriminate between a heavily obscured AGN scenario or an intrinsic X-ray weakness of these sources, we performed a radio (1.4GHz) stacking analysis, which did not reveal any detection and left open the questions about the origin of the X-ray weakness.
@@
Thanks to the successful launch of the James Webb Space Telescope (JWST; Gardner et al., 2023; Rigby et al., 2023), we are now able to investigate with high resolution and unprecedented sensitivity both the photometric and spectroscopic properties of galaxies up to (Carniani et al., 2024; Curtis-Lake et al., 2022; Robertson et al., 2022). Within this context, recent studies, exploiting both spectroscopic and imaging data from JWST, have revealed a large population of active galactic nuclei (AGN) at high redshift (Kocevski et al., 2023; Übler et al., 2023, 2024; Matthee et al., 2023; Maiolino et al., 2024b, 2023; Greene et al., 2023; Bogdán et al., 2023; Goulding et al., 2023; Kokorev et al., 2023; Furtak et al., 2023; Juodžbalis et al., 2024b; Scholtz et al., 2023a; Chisholm et al., 2024), providing the unique opportunity to study the properties of supermassive black hole (SMBH) and the AGN-galaxy coevolution, since very early times.
@@
Different works, selecting AGN at high- using JWST NIR and MIR photometry, showed that the AGN population at is larger than previously expected and that it is probably dominated by obscured, or even heavily obscured sources, as predicted by some coevolutional models Hopkins et al. (2008); Hickox & Alexander (2018). Yang et al. (2023a), taking advantage of the JWST-MIRI photometry of the Cosmic Evolution Early Release Science Survey (CEERS; Finkelstein et al., 2022), investigated the AGN population using spectral energy distribution (SED) modelling, and found a black-hole accretion rate density (BHARD) at dex higher than what was expected from previous X-ray AGN studies (Vito et al., 2016, 2018). Also Lyu et al. (2024) performing a similar analysis using the JWST/MIRI data of the SMILES survey selected a remarkable fraction of AGN among the MIRI detected sources () and found a statistically significant increase of the obscured AGN fraction with both AGN luminosity and redshift, as previously found in other works (Signorini et al., 2023; Gilli et al., 2022; Buchner et al., 2015; Aird et al., 2015).
@@
The existence of a larger-than-expected AGN population at was also shown by spectroscopic studies. For example, Maiolino et al. (2023) and Harikane et al. (2023), by selecting broad-line AGN (BLAGN, or TypeI) among JWST/NIRSpec spectra, found significant AGN excess at with respect to the AGN luminosity functions derived using X-ray data (Giallongo et al., 2019).
@@
JWST data not only allow the identification of a higher fraction of AGN at high- but also show some new and unexpected features in this early population of SMBHs. Studies have revealed that early SMBHs tend to be overmassive relative to the host galaxy stellar mass, when compared with the local AGN distribution (Maiolino et al., 2023; Bogdán et al., 2023; Furtak et al., 2023; Juodžbalis et al., 2024b), suggesting that the early stages of the coevolution between the SMBH and the host galaxy can be dominated by SMBH growth. This could further suggest that the preferred driving channel for the SMBH formation could be the so-called “Direct Collapse Black Hole” scenario, together with (or alternatively) episodes of super-Eddington accretion (Scholtz et al., 2023b).
@@
Another remarkable feature of high- newly discovered AGN is that they seem to be X-ray weak compared to the low-redshift AGN population. X-rays have been traditionally used to select AGN because at the typical AGN X-ray luminosities (), the level of contamination by stellar processes is low, allowing a generally pure AGN selection. However, even if X-ray photons are able to penetrate through dense environments, they are almost completely absorbed at Compton-thick (CTK) hydrogen column densities (), making the X-ray identification of heavily obscured AGN challenging (Vignali et al., 2015; Lanzuisi et al., 2018; Goulding et al., 2023; Maiolino et al., 2024a).
@@
Specifically, most of the newly selected AGNs, including BLAGN, lack any X-ray emission (Maiolino et al., 2024a; Yue et al., 2024), even if located in fields covered by some of the deepest extragalactic X-ray observations ever performed. Given the very few exceptions (Goulding et al., 2023; Kovacs et al., 2024; Maiolino et al., 2024a) of X-ray AGN detections in the early Universe, it is possible that the X-ray weakness could be due to intrinsic properties of high- AGN, as it was also suggested for some low- AGN (Simmonds et al., 2018; Zhang et al., 2023). In this view, AGN can be characterized by a different accretion-disk/coronal structure that can determine, for example, a larger ratio of the optical to X-ray emission () due to a much lower efficiency of the corona in producing X-ray photons, or simply lack of corona (Proga, 2005). Recently, Maiolino et al. (2024a), analysing the lack of X-ray emission in a large sample of high- sources unambiguously identified as AGN, suggested that their X-ray weakness could also be ascribed to the presence, in the inner region of the AGN, of a spherical distribution of clouds with CTK column densities and very low dust content, such as the Broad Line Region (BLR) clouds. In this hypothesis, high- BLAGN do not need to be characterized by any kind of intrinsic X-ray weakness, as their lack of X-ray emission would simply be caused by the inner CTK broad-line region gas.
@@
JWST spectroscopy can give us information on the nature of the observed sources, allowing to identify and investigate not only the population of BLAGN, which is probably only the tip of the iceberg of the entire AGN census but also the vast and hidden population of narrow-line AGN (NLAGN, also called TypeII). An additional significant difference between high- and low- AGN consists in the remarkably different physical environments constituted by their host galaxies. Indeed, early galaxies are systematically more metal-poor and characterized by younger stellar populations (Curti et al., 2023a) that increase their ionization parameters in the ISM (; Cameron et al., 2023; Curti et al., 2023a). These two effects have a significant impact on the selection of NLAGN, traditionally based on optical emission-line diagnostic diagrams, because the lower metallicities and the larger make star-forming galaxies (SFG) and AGN move towards each other and eventually overlap on the traditional diagnostic diagrams (Übler et al., 2023; Kocevski et al., 2023; Maiolino et al., 2023; Scholtz et al., 2023a; Mazzolari et al., 2024).
However, the identification of the elusive population of NLAGN is crucial in order to have a complete census of the AGN population, especially at high-, and to build up statistical AGN samples that can solve the problems on AGN identification and properties at high-.
@@
So far, the systematic search for NLAGN at has been attempted only by a few works. In Scholtz et al. (2023a) the authors selected 42 NLAGN among medium resolution (MR) spectra of the JADES survey (Eisenstein et al., 2023; Bunker et al., 2023). In particular, they used JADES spectra to select NLAGN through a series of rest-frame optical, rest-frame UV diagnostic diagrams, and by exploiting the detection of high-ionization emission lines, which usually require AGN as a photoionizing source. The selected sample allowed the authors to investigate the population of AGN up to , down to bolometric luminosities and host galaxy stellar masses of , vastly expanding the region of the parameter space populated by these objects at early times.
@@
In this work, we perform a similar analysis investigating the NLAGN population hidden among the sample of MR spectra collected in the CEERS survey. The paper is organized as follows. In Sec. 2, we present the data and the sample of spectra analyzed throughout this work and the techniques involved in the subsequent analysis: emission line fitting, spectral stacking, SED-fitting, and X-ray and radio stacking. In Sec. 3 we describe the different emission line diagnostic diagrams involved in the NLAGN selection and the results of the selection. Then in Sec. 4 we discuss the physical properties of the selected NLAGN: the AGN prevalence (Sec. 4.4), the distribution in velocity dispersion and obscuration (Sec. 4.5, Sec. 4.6), the bolometric luminosities (Sec. 4.7), the host galaxies properties (Sec. 4.8), and finally in Sec. 4.9 and Sec. 4.10 the X-ray and radio analysis, respectively.
We assume a flat CDM universe with , .
@@ 2.1 CEERS Observational Data
We use publicly available medium resolution (MR) JWST NIRSpec (Jakobsen et al., 2022) Micro-Shutter Assembly (MSA; Ferruit et al., 2022a) data from the CEERS program (Program ID:1345, Finkelstein et al., 2022). The CEERS NIRSpec observations consist of six pointings in the ‘Extended Groth Strip’ field (EGS; Rhodes et al., 2000; Davis et al., 2007), all of which utilized the three grating/filter combination of G140M/F100LP, G235M/F170LP, and G395M/F290LP. This provided a spectral resolution of R over the wavelength range of approximately 1–5 m (Jakobsen et al., 2022). Each grating/filter combination was observed for a total of 3107s for each pointing.
In particular, we considered the 313 spectra reduced and published by the CEERS collaboration in their latest data release DR-0.7111https://ceers.github.io/dr07.html. The data reduction was performed by the CEERS collaboration using the JWST Calibration Pipeline version 1.8.5 (Bushouse et al., 2022), using the CRDS context ”jwst_1029.pmap”. The spectra were reduced using processing standard pipeline parameters, with some specific deviation in particular regarding the ‘jump’ parameters and the so-called ”snowball” corrections. Nodded background subtraction was employed. The pipeline was instructed that all targets should be treated as point sources, determining each 1D spectrum to be extracted from the 2D spectral data file over a specified range of pixels in the cross-dispersion direction. Extraction apertures on the 2D spectra were determined for each object by interactive visual inspection and were specified explicitly for pipeline step ’extract_1d’. Most faint galaxies are compact, and the extraction apertures adopted for nearly all objects have heights ranging from 3 to 6 pixels (scale = 0.10 arcsec/pixel), with a median value of 4 pixels.
The default pipeline path-loss correction was employed. This calibration is based primarily on pre-flight modelling, assuming the targets are point sources and determining these corrections to be incomplete for significantly extended sources, but this is not the case for the large majority of the targets. Flux calibration uses the default reference files for the adopted CRDS context.
For this work, we used the spectra with the three single grating spectra combined together. The data from the individual MR gratings are resampled to a common wavelength vector in the overlapping regions, adopting the wavelength grating which in the overlap region has the worst spectral resolution. Flux values in the overlap regions are the average of the individual grating values weighted by the flux errors, excluding pixels affected by masked artefacts.
@@
Harikane et al. (2023) already investigated a limited sample of these spectra (i.e. only those at ) to look for BLAGN, selecting 10 AGNs at z = (2 of which are marked only as candidates) whose broad component is only seen in the permitted H or H lines and not in the forbidden [O iii] 5007 line. In the rest of the paper we mark the 8 most reliable BLAGN selected in Harikane et al. (2023) as the sample of high- BLAGN.
@@ 2.2 Emission line fitting
The NLAGN selection performed in this work is based on emission line diagnostics diagrams, where the AGN selection is provided by demarcation lines taken from the literature or by comparing the distribution of the sources with photoionization models, following the procedure done in Scholtz et al. (2023a).
We used a modified version of the publicly available code QubeSpec (Scholtz et al., 2023a; D’Eugenio et al., 2024) to fit the spectra of the sources. We fit the emission lines considering only small wavelength ranges around the single (or group of) emission lines (), with each emission line fitted using a single Gaussian component. The continuum was fit with a power law model, and in most of the spectra it was not detected. This approach for the fit of the continuum is sufficient for describing a narrow continuum range around an emission line of interest, and we found no instances of a strong continuum that required more sophisticated (e.g., stellar emission) modelling. QubeSpec uses a Bayesian approach, which requires defining prior probability distributions for each model parameter. In our fit, we assume log-uniform priors for the peak of the Gaussian describing each line and also for the continuum normalization. The prior on the lines full-width half maximum (FWHMs) is set to a uniform distribution spanning from the minimum resolution of the NIRSpec/MSA (km/s) up to a maximum of 700 km/s. For the fit of the high- BLAGN identified in Harikane et al. (2023) we also include in the fit a broad component in the H (with a FWHM uniform prior between km/s) to carefully disentangle the broad H emission from the narrow one, that is the only component used in the diagnostic diagrams.
The prior on the redshift is a normal distribution centred on the redshift obtained from the visual inspection with a standard deviation of 100 km/s.
To estimate the posterior probability distribution of all these parameters, we use a Markov chain Monte Carlo integrator (emcee; Foreman-Mackey et al., 2013).
@@
For the components of the [O iii] 4959,5008 and [N ii] 6548,84 doublets, we fixed the ratios between the two components to be 1:3 given by the relative Einstein coefficients of the two transitions.
Given the resolution of the NIRSpec MR spectra, we fit the [O ii] 3726,29, [Ne iv] 2422,24, C iv 1548,51, [O iii] 1660,66, C iii] 1907,10, and N iv] 1483,86 doublets as single Gaussians, centered at the mean wavelength between the two components.
The signal-to-noise ratio (S/N) of each line was evaluated considering the same equations reported in Mignoli et al. (2019) and Lenz & Ayres (1992), and requiring at least S/N for the detection. Then, we visually inspected all the detected lines to exclude any possible false detection. In case of undetection, we derived the upper limit to the flux of each line as 3 the error on the flux returned by QubeSpec.
While fitting the H line, we found 4 sources at for which the residuals from the narrow-line-only fit were too large, and a broad component was required. These sources were not selected as in Harikane et al. (2023) because in that work the authors investigated only sources at . We identify these sources as new BLAGN at low-; their spectra are presented in Appendix A.
In fitting all the BLAGN, we did not find any significant broad component either in the H or in the [O iii] 5007 lines.

### v3

# Narrow-line AGN selection in CEERS: Spectroscopic selection, physical properties, and X-ray and radio analysis

@@
Narrow-line AGN selection in CEERS: Spectroscopic selection, physical properties, and X-ray and radio analysis
The transformative era opened by the James Webb Space Telescope (JWST) on the high- Universe allows us to investigate the early stages of supermassive black hole (SMBH) evolution, with the first results showing a greater than expected number of active galactic nuclei (AGNs) at very early times. In this work, we spectroscopically select narrow-line AGNs (NLAGNs) among the publicly available medium-resolution spectra of the Cosmic Evolution Early Release Science Survey (CEERS). Using both traditional and newly identified emission line NLAGN diagnostics diagrams, we identified 52 NLAGNs at , on which we performed a detailed multiwavelength analysis. We also identified four new broad-line AGNs (BLAGNs), in addition to the eight previously reported BLAGNs. We found that the traditional BPT diagnostic diagrams are not suited to identifying high- AGNs, while most of the high- NLAGN were selected using the recently proposed AGN diagnostic diagrams based on the [O iii] 4363 auroral line or high-ionization emission lines.We compared the emission line velocity dispersion and the obscuration levels of the sample of NLAGNs with those of the parent sample without finding significant differences between the two distributions, suggesting a population of AGNs heavily buried and not significantly impacting the host galaxies’ physical properties, as was further confirmed by spectral energy distribution fitting. The bolometric luminosities of the high- NLAGNs selected in this work are dex below the ones sampled by surveys before JWST, potentially explaining the weak impact of these AGNs. Finally, we investigated the X-ray properties of the selected NLAGNs and of the sample of high- BLAGNs. We found that all but four NLAGNs are undetected in the deep X-ray image of the field, as well as all the high- BLAGNs. We did not obtain a detection even by stacking the undetected sources, resulting in an X-ray weakness of dex from what was expected based on their bolometric luminosities. To discriminate between a heavily obscured AGN scenario or an intrinsic X-ray weakness of these sources, we performed a radio (1.4GHz) stacking analysis, which did not reveal any detection and left open the questions about the origin of the X-ray weakness.
@@
Thanks to the successful launch of the James Webb Space Telescope (JWST; Gardner23; Rigby23), we are now able to investigate with a high resolution and an unprecedented sensitivity both the photometric and spectroscopic properties of galaxies up to (Carniani24; CurtisLake22; Robertson22). Within this context, recent studies, exploiting both spectroscopic and imaging data from JWST, have revealed a large population of active galactic nuclei (AGNs) at high redshift (Kocevski23; ubler23; ubler23b; Matthee23; Maiolino23a; Maiolino23c; Greene23; Bogdan23; Goulding23; Kokorev23; Furtak23; Juodzbalis24; Scholtz23b; Chisholm24), providing a unique opportunity to study the properties of supermassive black hole (SMBH) and AGN-galaxy coevolution since very early times.
@@
Different works, selecting AGNs at high- using JWST NIR and MIR photometry, have shown that the AGN population at is larger than previously expected and that it is probably dominated by obscured or even heavily obscured sources, as is predicted by some coevolutional models Hopkins08; hickox18. Yang23, taking advantage of the JWST-MIRI photometry of the Cosmic Evolution Early Release Science Survey (CEERS; Finkelstein22), investigated the AGN population using spectral energy distribution (SED) modeling and found a black-hole accretion rate density (BHARD) at dex higher than what was expected from previous X-ray AGN studies (vito16; vito18). Also, Lyu23, performing a similar analysis using the JWST/MIRI data of the SMILES survey, selected a remarkable fraction of AGNs among the MIRI-detected sources () and found a statistically significant increase in the obscured AGN fraction with redshift, as had previously been found in other works (Signorini23; gilli22; buchner15; aird15).
@@
The existence of a larger-than-expected AGN population at has also been shown by spectroscopic studies. For example, Maiolino23c and Harikane23, by selecting broad-line AGNs (BLAGNs, or type I) among JWST/NIRSpec spectra, found a significant AGN excess at with respect to the AGN luminosity functions derived using X-ray data (giallongo19).
@@
JWST data not only allow for the identification of a higher fraction of AGNs at high- but also show some new and unexpected features in this early population of SMBHs. Studies have revealed that early SMBHs tend to be overmassive relative to the host galaxy stellar mass, when compared with the local AGN distribution (Maiolino23c; Bogdan23; Furtak23; Juodzbalis24), suggesting that the early stages of the coevolution between the SMBH and the host galaxy can be dominated by SMBH growth. This could further suggest that the preferred driving channel for the SMBH formation is the so-called “direct collapse black hole” scenario, together with (or alternatively) episodes of super-Eddington accretion (Scholtz_2023_GN-z11).
@@
Another remarkable feature of high- newly discovered AGNs is that they seem to be X-ray-weak compared to the low-redshift AGN population. X-rays have traditionally been used to select AGNs because at typical AGN X-ray luminosities (), the level of contamination by stellar processes is low, allowing for a generally pure AGN selection. However, even if X-ray photons are able to penetrate through dense environments, they are almost completely absorbed at Compton-thick (CTK) hydrogen column densities (), making the X-ray identification of heavily obscured AGNs challenging (vignali15; lanzuisi18; Goulding23; Maiolino24_X).
@@
Specifically, most of the newly selected AGNs, including BLAGNs, lack any X-ray emission (Maiolino24_X; Yue24), even if they are located in fields covered by some of the deepest extragalactic X-ray observations ever performed. Given the very few exceptions (Goulding23; Kovacs24; Maiolino24_X) of X-ray AGN detections in the early Universe, it is possible that the X-ray weakness could be due to intrinsic properties of high- AGNs, as has also been suggested for some low- AGNs (simmonds18; Zhang23). In this view, AGNs can be characterized by a different accretion-disk and/or X-ray corona structure that can determine, for example, a larger ratio of the optical to X-ray emission () due to a much lower efficiency of the corona in producing X-ray photons, or simply a lack of corona (Proga05). Recently, Maiolino24_X, analyzing the lack of X-ray emission in a large sample of high- sources unambiguously identified as AGNs, suggested that their X-ray weakness could also be ascribed to the presence, in the inner region of the AGNs, of a spherical distribution of clouds with CTK column densities and very low dust content, such as the broad-line region (BLR) clouds. In this hypothesis, high- BLAGNs do not need to be characterized by any kind of intrinsic X-ray weakness, as their lack of X-ray emission would simply be caused by the inner CTK BLR gas.
@@
JWST spectroscopy can give us information on the nature of the observed sources, allowing us to identify and investigate not only the population of BLAGNs, which is probably only the tip of the iceberg of the entire AGN census, but also the vast and hidden population of narrow-line AGNs (NLAGNs, also called type II). An additional significant difference between high- and low- AGNs lies in the remarkably different physical environments constituted by their host galaxies. Indeed, early galaxies are systematically more metal-poor and characterized by younger stellar populations (Curti23) that increase their ionization parameters in the ISM (; Cameron23; Curti23). These two effects have a significant impact on the selection of NLAGNs, traditionally based on optical emission-line diagnostic diagrams, because the lower metallicities and the larger make star-forming galaxies (SFGs) and AGNs move toward each other and eventually overlap on the traditional diagnostic diagrams (ubler23; Kocevski23; Maiolino23c; Scholtz23b; Mazzolari24b).
@@
However, the identification of the elusive population of NLAGNs is crucial in order to have a complete census of the AGN population, especially at high-, and to build up statistical AGN samples that can solve the problems in AGN identification and properties at high-. So far, a systematic search for NLAGNs at has only been attempted by a few works. In Scholtz23b, the authors selected 42 NLAGNs among medium-resolution (MR) spectra of the JADES survey (Eisenstein23; bunker23b). In particular, they used JADES spectra to select NLAGN through a series of rest-frame optical, rest-frame UV diagnostic diagrams, and by exploiting the detection of high-ionization emission lines, which usually require an AGN as a photoionizing source. The selected sample allowed the authors to investigate the population of AGNs up to , down to bolometric luminosities of and host galaxy stellar masses of , vastly expanding the region of the parameter space populated by these objects at early times.
@@
In this work, we perform a similar analysis, investigating the NLAGN population hidden among the sample of MR spectra collected in the CEERS survey. The paper is organized as follows. In Sec. 2, we present the data and the sample of spectra analyzed throughout this work and the techniques involved in the subsequent analysis: emission line fitting, spectral stacking, SED-fitting, and X-ray and radio stacking. In Sec. 3, we describe the different emission line diagnostic diagrams involved in the NLAGN selection and the results of the selection. Then, in Sec. 4, we discuss the physical properties of the selected NLAGNs: the AGN prevalence (Sec. 4.4), the distribution in velocity dispersion and obscuration (Sec. 4.5, Sec. 4.6), the bolometric luminosities (Sec. 4.7), the host galaxies properties (Sec. 4.8), and finally, in Sec. 4.9 and Sec. 4.10, the X-ray and radio analysis, respectively. We assume a flat CDM Universe with , , and .
@@ 2.1 CEERS observational data
We used publicly available MR JWST NIRSpec (Jakobsen22) Micro-Shutter Assembly (MSA; Ferruit23) data from the CEERS program (Program ID:1345, Finkelstein22). The CEERS NIRSpec observations consist of six pointings in the ‘Extended Groth Strip’ field (EGS; Rhodes00; Davis07), all of which utilized the three grating and filter combination of G140M/F100LP, G235M/F170LP, and G395M/F290LP. This provided a spectral resolution of R over the wavelength range of approximately 1–5 m (Jakobsen22). Each grating and filter combination was observed for a total of 3107s for each pointing.
In particular, we considered the 313 spectra reduced and published by the CEERS collaboration in their latest data release, DR-0.7111https://ceers.github.io/dr07.html. The data reduction was performed by the CEERS collaboration using the JWST Calibration Pipeline version 1.8.5 (Bushouse22), using the CRDS context “jwst_1029.pmap.” The spectra were reduced using standard processing pipeline parameters, with some specific deviation in particular regarding the “jump” parameters and the so-called “snowball” corrections. Nodded background subtraction was employed. The pipeline was instructed that all targets should be treated as point sources, determining each 1D spectrum to be extracted from the 2D spectral data file over a specified range of pixels in the cross-dispersion direction. Extraction apertures on the 2D spectra were determined for each object by interactive visual inspection and were specified explicitly for the pipeline step “extract_1d.” Most faint galaxies are compact, and the extraction apertures adopted for nearly all objects have heights ranging from 3 to 6 pixels (scale = 0.10 arcsec/pixel), with a median value of 4 pixels.
The default pipeline path-loss correction was employed. This calibration is based primarily on pre-flight modeling, assuming the targets are point sources and determining these corrections to be incomplete for significantly extended sources, but this is not the case for the large majority of the targets. Flux calibration uses the default reference files for the adopted CRDS context.
For this work, we used the spectra with the three single grating spectra combined together. The data from the individual MR gratings were resampled to a common wavelength vector in the overlapping regions, adopting the wavelength grating that in the overlap region has the worst spectral resolution. The flux values in the overlap regions are the average of the individual grating values weighted by the flux errors, excluding pixels affected by masked artefacts.
@@
Harikane23 already investigated a limited sample of these spectra (i.e., only those at ) to look for BLAGN, selecting 10 AGNs at z = (2 of which are marked only as candidates) whose broad component is only seen in the permitted H or H lines and not in the forbidden [O iii] 5007 line. In the rest of the paper we mark the 8 most reliable BLAGNs selected in Harikane23 as the sample of high- BLAGNs.
@@ 2.2 Emission line fitting
The NLAGN selection performed in this work is based on emission line diagnostics diagrams, where the AGN selection is provided by demarcation lines taken from the literature or by comparing the distribution of the sources with photoionization models, following the procedure done in Scholtz23b222https://github.com/honzascholtz/Qubespec.
We used a modified version of the publicly available code QubeSpec (Scholtz23b; DEugenio24) to fit the spectra of the sources. We fit the emission lines considering only small wavelength ranges around the single (or group of) emission lines (), with each emission line fit using a single Gaussian component. The continuum was fit with a power law model, and in most of the spectra it was not detected. This approach for the fit of the continuum is sufficient for describing a narrow continuum range around an emission line of interest, and we found no instances of a strong continuum that required more sophisticated (e.g., stellar emission) modeling. QubeSpec uses a Bayesian approach, which requires defining prior probability distributions for each model parameter. In our fit, we assume log-uniform priors for the peak of the Gaussian describing each line and also for the continuum normalization. The prior on the lines full-width half maximum (FWHMs) is set to a uniform distribution spanning from the minimum resolution of the NIRSpec/MSA (km/s) up to a maximum of 700 km/s. For the fit of the high- BLAGNs identified in Harikane23 we also include in the fit a broad component in the H (with a FWHM uniform prior between km/s) to carefully disentangle the broad H emission from the narrow one, which is the only component used in the diagnostic diagrams.
The prior on the redshift is a normal distribution centered on the redshift obtained from the visual inspection with a standard deviation of 100 km/s.
To estimate the posterior probability distribution of all these parameters, we use a Markov chain Monte Carlo integrator (emcee; ForemanMackey13).
@@
For the components of the [O iii] 4959,5008 and [N ii] 6548,84 doublets, we fixed the ratios between the two components to be 1:3 given by the relative Einstein coefficients of the two transitions. Given the resolution of the NIRSpec MR spectra, we fit the [O ii] 3726,29, [Ne iv] 2422,24, C iv 1548,51, [O iii] 1660,66, C iii] 1907,10, and N iv] 1483,86 doublets as single Gaussians, centered at the mean wavelength between the two components. The signal-to-noise ratio (S/N) of each line was evaluated considering the same equations reported in mignoli19 and Lenz92:
where is the sampling interval of the spectrum, is the peak S/N (i.e., , where are the flux density and the noise level at the central pixel), and C is a proportionality constant whose values are reported in Table 1 of Lenz92. We also checked the S/N using the following equation:
where and are the flux densities and associated errors for flux bins falling inside the Gaussian profile of the fit, and we found results consistent with those derived from Eq. 1.
We required at least S/N for the detection. Then, we visually inspected all the detected lines to exclude any possible false detection. In case of undetection, we derived the upper limit to the flux of each line as 3 errors (with the error returned by QubeSpec).
While fitting the H line, we found 4 sources at for which the residuals from the narrow-line-only fit were too large, and a broad component was required. These sources were not selected as in Harikane23 because in that work the authors investigated only sources at . We identify these sources as new BLAGNs at low-; their spectra are presented in Appendix A.
In fitting all the BLAGNs, we did not find any significant broad component either in the H or in the [O iii] 5007 lines.